The perimeter of a rectangle is 220 inches. THe lenght exceeds the width by 40 Inches. What is the lenght and the width?
Answer by MathTherapy(10553)

Let the width of the rectangle be w
Then the length of the rectangle is w + 40
Then, 2w + 2(w + 40) = 220
2w + 2w + 80 = 220
4w = 140
w = 35
Therefore, width = 35, and length = 75 (35 + 40).
Check:
220 (perimeter) = 2(35) + 2(35 + 40)
220 = 70 + 150 (True)